Contents

Operating Systems and Systems Integration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 1 Is this a theoretical subject? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 2 NO! This is practical! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 3 Mostly based on Linux. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 4 Nick Loves Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 5 So why does Nick love Linux? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 6 More reasons why Nick Loves Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 7 Open Standards and Protocols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 8 Standards that are not so open . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 9 Embrace and Extend. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 10 Other “Industry Standards” . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 11 What is “Systems Integration”? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 12 Why From Many Companies? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 13 Monoculture (One supplier). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 14 How Best to Integrate? Standards . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 15 So what answers are there? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 16 Samba — Systems Integration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 17 Practical: 60 hours of laboratory workshops! . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Slide 18 Introducton to OSSI - p. 1/18

Operating Systems and Systems Integration

Is this a boring subject only about the theory of the internal operation of an operating system???

Introducton to OSSI - p. 2/18

Is this a theoretical subject?

Hmmm, operating systems: sounds like a theoretical subject At university, a lot of theory from a text book Are we going to spend lots of time copying from a text book? …

Introducton to OSSI - p. 3/18

NO! This is practical!

This subject aims to provide you with practical skills that you will find useful in your workplace. You learn here by doing. I even understand this subject too!

Introducton to OSSI - p. 4/18

Mostly based on Linux

… with some comparisons with Windows 2000. Aim to support career certification, e.g., Red Hat Certified Engineer, Linux Professional Institute

OSSI — ver. 1.1  Introducton to OSSI - p. 5/18

Introducton to OSSI - p. 6/18

So why does Nick love Linux?

It is free software Free as in freedom … also, free as in free beer It works really well It can “glue” many other things together Like Lego; can build anything I want

Introducton to OSSI - p. 7/18

More reasons why Nick Loves Linux

Built on cooperation Great for Internet applications Only fully compliant tcp^{/}ip We can see and modify the source code to any of it that we want to Based on standards Uses open protocols

Introducton to OSSI - p. 8/18

Open Standards and Protocols

What do I mean by “Open Standard?” Can freely download the standard, not pay a huge fee just to read it Agreed to by open discussion: barrier to participation is lack of ability, not lack of money! Examples of open standards: tcp^{/}ip, http, ssh, ftp, dns, dhcp, Perl, ldap, smtp, tls, many, many more

Introducton to OSSI - p. 9/18

Standards that are not so open

Many standards are developed in a less open way Example: the wep protocol, used to provide “privacy” for wireless lans Developed behind closed doors, announced to the public Cracked almost immediately: a bad design Not open to inspection and peer review

Introducton to OSSI - p. 10/18

Embrace and Extend

Many companies use “standards” But change them a little bit to give a competitive advantage Example: Kerberos

Developed at MIT as open source software,

Improved by programmers all round the world

Used by Microsoft for authentication in Active Directory Microsoft changed the interpretation of one small part of the protocol Effect: all Kerberos clients can use a Microsoft Kerberos server But Microsoft clients will refuse to work with anything except a Microsoft server.

Introducton to OSSI - p. 11/18

Other “Industry Standards”

Consider the Microsoft Office suite The layout is a secret OpenOffice.org has developed an office suite that can read and write Office documents Very hard work:

a moving target

Much effort by ms to make them very hard to read, and even harder to write

Introducton to OSSI - p. 12/18

What is “Systems Integration”?

It involves combining products from many companies into a system Other words: interoperation, compatibility Very important: it’s not enough just to learn one product

Introducton to OSSI - p. 13/18

Why From Many Companies?

Why not just from one supplier?

Introducton to OSSI - p. 14/18

Monoculture (One supplier)

Advantages: Fewer system integration problems Less skill required All training from the one vendor Disadvantages: Vulnerability in one is a vulnerability in all (e.g., So.Big, Blaster, SirCam, NIMBDA worms, Outlook viruses) One supplier cannot make everything: E.g., Cisco sell more routers and switches than Microsoft

Introducton to OSSI - p. 15/18

How Best to Integrate? Standards

Using Open Standards that are free of patents and other restrictions on use … But every vendor says they support the standards! Some standards are more open than others!

Introducton to OSSI - p. 16/18

So what answers are there?

Free software supports open standards So open, you can read the source code and see how it works! You can even change it to suit your needs, and if you find problems, you can fix them! Aim for interoperation, not for exclusion or market benefit

Introducton to OSSI - p. 17/18

Samba — Systems Integration

Samba allows a Linux or Unix or Macintosh machine to talk with Windows and the other way round Free Software A Linux machine can be an NT-4 compatible Primary Domain Controller Can also be an AD member Very stable, high quality Used by many companies to interoperate with Windows

Introducton to OSSI - p. 18/18

Practical: 60 hours of laboratory workshops!

We start this week! So what will we do? In the first class, we partition hard disks … then install Linux. See you there!